OpenStudy (anonymous):

what is the positive and a negative of |x|>=-1

OpenStudy (anonymous):

please i need help

OpenStudy (jdoe0001):

\(\bf |x|\ge-1\implies \begin{cases} +(x)\ge-1\implies x \ge -1\\ \quad \\ \bf -(x)\ge-1\implies x \le 1 \end{cases}\implies -1\le x\le 1\) all absolute value expressions have 2 scenarios, thus 2 values for "x"